# Forecasting Response to Treatment with Global Deep Learning and Patient-Specific Pharmacokinetic Priors

### v6

@@ Abstract
Forecasting healthcare time series is crucial for early detection of adverse outcomes and for patient monitoring. Forecasting, however, can be difficult in practice due to noisy and intermittent data. The challenges are often exacerbated by change points induced via extrinsic factors, such as the administration of medication. To address these challenges, we propose a novel hybrid global-local architecture and a pharmacokinetic encoder that informs deep learning models of patient-specific treatment effects. We showcase the efficacy of our approach in achieving significant accuracy gains for a blood glucose forecasting task using both realistically simulated and real-world data. Our global-local architecture improves over patient-specific models by 9.2-14.6%. Additionally, our pharmacokinetic encoder improves over alternative encoding techniques by 4.4% on simulated data and 2.1% on real-world data. The proposed approach can have multiple beneficial applications in clinical practice, such as issuing early warnings about unexpected treatment responses, or helping to characterize patient-specific treatment effects in terms of drug absorption and elimination characteristics.
@@
Multi-Input Forecasting, Glucose, Pharmacokinetics, Deep Learning
@@ 1 Introduction
Forecasting healthcare time series is crucial for early detection of adverse outcomes and patient monitoring (Churpek et al., 2016; Gerry et al., 2020). Notably, healthcare time series data are often affected by exogenous factors, such as medication, exercise, or diet. Accurate forecasting systems should strive to harness the predictive potential of these factors. While the inclusion of exogenous variables can significantly enhance prediction accuracy (Olivares et al., 2022a), there are instances in which their integration yields minimal to no improvement (McShinsky and Marshall, 2020; Rubin-Falcone et al., 2020).
Figure 1: Insulin dose is usually known and recorded in data, whereas plasma insulin concentration is unobservable and subject to changes over time.
Leveraging the predictive power of exogenous variables presents unique challenges, especially in healthcare, where patient signals and exogenous factors, such as medication dose, often exhibit a temporal resolution mismatch. Furthermore, while the medication dose is usually known, we cannot directly observe the drug absorption and elimination parameters or plasma concentration of the drug, as illustrated in Fig. 1. Consequently, their impact on the predictive signal remains hidden.
We consider the blood glucose forecasting task while taking into account hidden medication effects. In our data, blood glucose measurements are taken every five minutes, and exogenous variables such as carbohydrate ingestion or application of insulin are recorded as instantaneous points in time. To model hidden temporal effects of such exogenous events, we propose to modify contemporary deep learning forecasting models by adding a pharmacokinetic encoder. The encoder leverages pharamacokinetic knowledge to generate patient-specific plasma insulin concentration profiles. In our framework, global parameters of a forecasting architecture are shared across multiple patients (Semenoglou et al., 2021), but the parameters of the encoder are learned from the individual patent’s historical data to address inter-subject variability in response to treatment.
@@
  • Pharmacokinetic Encoder. We introduce a novel architecture module that generates plasma drug concentration profiles to capture time-dependent medication treatment effects.
@@
  • Hybrid Global-Local Architecture. We bolster global architectures (with parameters shared across patients) with a learnable patient-specific pharmacokinetic encoder.
@@
  • State-of-the-Art Results. We obtain significant accuracy improvements on a large scale simulated and real-world blood glucose forecasting datasets.
1.1 Glucose Forecast Background
Type I diabetes (T1DM) patients must regularly monitor and regulate their blood glucose values to prevent hypoglycemic and hyperglycemic events, which can drastically impact organs or even be fatal. Insulin dosages, T1DM’s most crucial treatment, are represented as instantaneous events in data, which can hide their long-lasting effects. Effectively capturing the dynamic effects of medication over time is crucial to accurately predict the impact of insulin on blood glucose.
Pharmacokinetic modeling is a natural solution to the aforementioned problem, as it can unveil medication effects through knowledge of drug movement through the body. For example, pharmacokinetics provides a well-established understanding of plasma insulin concentration changes over time based on drug absorption and elimination characteristics  (Binder et al., 1984).
For extravascular injections, the plasma concentration-time profile exhibits right-skewed characteristics (Rosenbaum, 2017). Fig. 2 depicts pharmacokinetic modeling of plasma insulin concentration over time for different bolus and basal insulin types. Generally, bolus doses are quick-acting and, thus, often taken before meals compared to basal insulin, which is longer-acting and serves to regulate blood glucose levels throughout the day.
In Fig. 2, the area under the concentration curve measures drug bioavailability, the fraction of the absorbed dose that reaches the intended site of systemic circulation intact. Bioavailability can vary based on many factors, including type of medication, dosage, and patient-specific characteristics. For instance, the bioavailability of an intravenous dose of any drug is close to 100 percent (Søeborg et al., 2012). However, for subcutaneous injections, bioavailability is lower due to unintended drug absorption or elimination.
@@ 1.1 Glucose Forecast Background
Figure 2: Pharmacokinetic modeling of plasma insulin concentration over time.
1.2 Previous Work
Neural Forecasting. In recent years, neural networks have demonstrated significant improvements compared to traditional forecasting algorithms, following the adoption of the cross-learning technique (Smyl, 2019; Semenoglou et al., 2021), which involves utilizing expressive global models to leverage data from vast collections of related time series.
The success of this global modeling approach has seen deep forecasting methods become ubiquitous in industrial forecasting systems (Wen et al., 2017; Lim et al., 2021), and in academia (Oreshkin et al., 2020).
Blood Glucose Forecasting. Neural Forecasting literature has permeated into the field of blood glucose level prediction, outperforming the accuracy of classic statistical methods (Xie and Wang, 2020). Past research includes various neural network designs ranging from Multi Layer Perceptrons (MLP; Jahangir et al. 2017), Recurrent Neural Networks (RNN; Fox et al. 2018; Rabby et al. 2021; Rubin-Falcone et al. 2022) to Temporal Convolutional Networks (TCN; Li et al. 2019).
Despite recent advancements in blood glucose forecasting, two significant challenges remain. Firstly, existing architectures struggle to effectively utilize sparse exogenous variables with the exception of Rubin-Falcone et al. (2022). Medication and carbohydrate values are recorded as sparse values in data, which may contribute to model difficulty in learning the effects of treatment on blood glucose levels. Early solutions propose summing up exogenous values to the current point within an input window to replace sparse features as shown in Fig. 3. However, this “sum-total” preprocessing approach fails to capture time-dependent changes in medication, which can be achieved through pharmacokinetic modeling of plasma insulin concentration. Despite this, deep learning methods used to forecast blood glucose do not do not consider pharmacokinetics to inform model training. Secondly, many studies (Xie and Wang, 2020; Rubin-Falcone et al., 2022; Zaidi et al., 2021) rely on individualized networks for each patient, with limited adoption of cross-learning approaches that could enhance forecasting models by leveraging shared knowledge across patients. Our proposed approach address both challenges in a novel hybrid global-local modeling approach that learns patient-specific pharmacokinetics to improve deep learning model capabilities in forecasting blood glucose levels.

### v7

# Forecasting Treatment Response with Deep Pharmacokinetic Encoders

@@ Abstract
Forecasting healthcare time series data is vital for early detection of adverse outcomes and patient monitoring. However, forecasting is challenging in practice due to variable medication administration and unique pharmacokinetic (PK) properties for each patient. To address these challenges, we propose a novel hybrid global-local architecture and a PK encoder that informs deep learning models of patient-specific treatment effects. We showcase the efficacy of our approach in achieving significant accuracy gains for a blood glucose forecasting task using both realistically simulated and real-world data. Our hybrid global-local architecture improves over patient-specific models by 15.8%  on average. Additionally, our PK encoder surpasses baselines by up to 16.4% on simulated data and 4.9% on real-world data for individual patients during critical events of severely high and low glucose levels. Our code is available at https://github.com/PotosnakW/neuralforecast/tree/pk_paper_code.
@@
Multi-Input Forecasting, Healthcare, Pharmacokinetics, Deep Learning
@@ 1 Introduction
Leveraging the predictive power of exogenous variables presents unique challenges, especially in healthcare, where the frequency of patient signals and exogenous factors, such as medication doses or treatments, often exhibit a mismatch in temporal resolution. Furthermore, while the medication dose is usually known, one cannot directly observe the plasma concentration of the drug, as illustrated in Fig. 1. Consequently, the impact of dynamic changes in treatment resulting from drug absorption and elimination remains hidden.
Figure 1: Insulin dose is usually known and recorded in data, whereas plasma insulin concentration is unobservable and varies over time.
We consider the blood glucose forecasting task while taking into account latent, time-dependent treatment effects. Patients with diabetes must regularly monitor and regulate their blood glucose values to prevent hypoglycemic and hyperglycemic events, which can cause irreversible organ damage or, in severe cases, be fatal. Hyperglycemic and hypoglycemic episodes are responsible for numerous hospital admissions each year. A 2009 report on hospital admissions for hyperglycemia indicates over 500,000 patient-days and costs of \$2.4 billion each year for hyperglycemia alone (Kitabchi et al., 2009). Another report on hypoglycemic hospital admissions suggests a stable rate of 5.6-7.3 admissions per 1,000 patients annually between 2016-2019 (Pilla et al., 2021). With over 38 million patients with Diabetes Mellitus in the United States alone, this equates to a conservative estimate of 212,800 patients hospitalized for hypoglycemic episodes each year. Capturing the dynamic effects of insulin is crucial for accurately predicting its impact on blood glucose levels to anticipate and proactively address abnormal levels to prevent adverse patient events.
@@
  • Hybrid Global-Local Architecture. We bolster global architectures (with parameters shared across patients) with a learnable patient-specific encoder.
  • Pharmacokinetic (PK) Encoder. We introduce a novel deep learning model-agnostic module that generates plasma concentration profiles to capture time-dependent treatment effects.
  • Multiple Dose Effects Framework. We provide a computationally efficient framework for encoding the effects of multiple doses. We verify this approach mathematically under the assumption of linear PK.
@@
  • State-of-the-Art Results. We obtain significant accuracy improvements on large scale simulated and real-world patient blood glucose forecasting datasets.
1.1 Pharmacokinetics Background
PK modeling is a natural solution to the aforementioned problem, as it can unveil medication effects through knowledge of drug movement within the body. For example, PK provides a well-established understanding of plasma insulin concentration changes over time based on drug absorption and elimination characteristics  (Binder et al., 1984). Empirical and invasive studies show that concentration-time profiles for subcutaneous injections follow right-skewed curves. The degree of skew varies with insulin type and patient-specific factors, as shown in Fig. 2, where the area under the concentration curve measures drug bioavailability, or the fraction of the absorbed dose that reaches the intended site of systemic circulation intact. Most drugs used in clinical practice are assumed to follow linear PK (Rosenbaum, 2017). Under this model, the area under the concentration curve is directly proportional to the drug dose. A complete reference for supporting PK information regarding concentration-time profiles, insulin analogs, bioavailability, first-order kinetics, and linear PK can be found in Appendix B.
@@ 1.1 Pharmacokinetics Background
Figure 2: Pharmacokinetic modeling of plasma insulin concentration over time for various insulin types.
2 Related Work
Blood Glucose Forecasting. Neural Forecasting has previously been applied to the field of blood glucose forecasting, outperforming the accuracy of classic statistical methods. RNNs/LSTMs have been employed both as models and baselines in numerous studies (Qian et al., 2021; Li et al., 2019a; Fox et al., 2018; Rabby et al., 2021; Rubin-Falcone et al., 2020, 2022; Miller et al., 2020). Similarly, CNNs (Xie and Wang, 2020; Li et al., 2019a, b) and MLPs (McShinsky and Marshall, 2020; Miller et al., 2020; Jahangir et al., 2017) have been utilized in proposed models and as baselines.
Blood Glucose Treatment Effect Modeling. Previous studies have successfully modeled treatment effects on patient blood glucose levels. Miller et al. present one of the first hybrid statistical-and-physiological model that combines glucose-insulin dynamics and deep learning (Miller et al., 2020). They present the Deep T1D Simulator (DTD-Sim) model, which uses a deep state-space model to describe parameters representing insulin and carbohydrate absorption rates and sensitivities of the UVA/Padova simulator that simulates blood glucose levels. Odnoblyudova et al. demonstrated enhanced accuracy in predicting blood glucose levels using Multi-output Gaussian Processes (MOGPs) to model composite effects of carbohydrate and fat components from meals. (Odnoblyudova et al., 2023).
Despite recent advancements in forecasting treatment effects, significant challenges remain, including:
Sparse data.
Limited treatment data can hinder deep learning models from accurately learning treatment responses (Rubin-Falcone et al., 2022; Hussain et al., 2021). To address this challenge, we integrate domain knowledge into the deep learning models, allowing them to more effectively encode the time-dependent effects of these variables and improve forecasting performance.
Composite treatment effects.
Unlike meals that consist of various components such as carbohydrates and fats, insulin is a homogenous protein hormone composed of a single type of molecule. Therefore, modeling its component effects is not applicable (Odnoblyudova et al., 2023). Instead, we aim to model patient-specific holistic effects of the drug as well as composite effects of multiple doses.
Inflexible models.
Previous glucose forecasting models, like the UVA/Padova simulator, are specific to blood glucose dynamics and inflexible to other forecasting targets and variables (Visentin et al., 2016; Miller et al., 2020). Our approach leverages a deep learning model to generate predictions. This distinction provides greater flexibility, as our forecasting model is not constrained by a specific physiological model and can accommodate various deep learning models and forecasting targets and inputs.
Model application scope.
Prior work that integrates PK domain knowledge in deep learning models has been used in the context of predicting disease progression for cancer and COVID-19 with long-term horizons of interest (Hussain et al., 2021; Qian et al., 2021). In contrast, we leverage linear PK assumptions to create a framework that can be directly applied to various medications used in clinical practice and is easily adapted to other treatments.
Individualized networks.
Studies often use individualized networks, limiting cross-learning. Our novel global-local architecture allows efficient training of a more accurate global model applicable to multiple patients.
3 Methods
We aim to forecast blood glucose levels 30 minutes into the future consistent with prior work (Rubin-Falcone et al., 2022; Xie and Wang, 2020; Li et al., 2019b), based on a 10 hour history of blood glucose, carbohydrate (CHO), insulin basal, and insulin bolus values. We choose a 10-hour duration to balance sufficient historical information with GPU memory constraints. the duration of the rapid-acting insulin (Humalog and Novolog) used in insulin pumps by patients in our cohort typically ranges from 2 to 5 hours (ins, 2018, ). Therefore, the selection of a 10-hour lag time can fully cover multiple dose durations, allowing the model to learn treatment effects across various dosages and durations.
3.1 PK Encoder
Medication administrations are represented as instantaneous, sparse variables in time series data. We propose a novel PK encoder that informs deep learning models of time-dependent plasma drug concentration to enable more accurate forecasting of treatment outcomes as shown in Fig. 3. The encoder consists of a function, $C$, that takes in sparse exogenous variables and generates a concentration-time profile for a specified dose, $\textbf{x}^{(i)}$, using the following equation:
$C(t,\textbf{x}^{(i)},\textbf{k}^{(i)})=\frac{\textbf{x}^{(i)}}{\textbf{k}^{(i)}\sqrt{2\pi}t}\exp\{-\frac{1}{2{(\textbf{k}^{(i)})}^{2}}(\log(t)-1)^{2}\}$ (1)
where $t$ is the time and $\textbf{k}^{(i)}$ is a PK parameter affecting elimination rate that specific to patient, $i$. Here, $C$ is equivalent to the log-normal probability density function (pdf) with $\sigma=\textbf{k}$ and $\mu=1$ and a scaling factor, x. The intuition for the log-normal shape of the concentration curves stems from pharmacokinetics literature as outlined in Section 1.1 and Appendix B.
Figure 3: Medication administrations are represented as sparse variables in time series data. We propose a pharmacokinetic (PK) encoder to effectively capture time-dependent plasma drug concentration (Conc.).
Under linear PK assumptions, PK parameters, including bioavailability, are assumed to be constant with respect to the drug. Given this assumption, we show mathematically that the area under the concentration curve is bounded by the dose in Proposition A.1 in Appendix A.5. To model bioavailability and incorporate dose information directly into the concentration curves, we use the dose, x, as a scaling factor in $C$. This ensures that the area under the curve integrates to the dose value, under the simplifying assumption that all insulin administrations adhere to 100% bioavailability, which we show is upper bounded by the dose. Scaling the concentration curve changes its magnitude but does not alter the PK parameters or shape of the curve.
Let $N$ be the number of subjects in the data. The PK parameter $\textbf{k}\in[0,1]^{N\times 1}$ that controls the shape of the concentration curve, is represented as network embedding weight that is initialized between values of 1 and 2 prior to model training. The initial k values are tuned as model hyperparameters.
The encoder is trained end-to-end with the model. During model training, sparse insulin features are passed into the PK encoder along with their respective $\textbf{k}^{(i)}$ parameters, as shown in Fig. 4. The encoder than replaces sparse medication features in the training window with the corresponding concentration curve. For a single dose, x, of insulin administered at time, $t$, we obtain the concentration, $\widecheck{\textbf{x}}^{(i)}$, at that time for patient, $i$:
$\widecheck{\textbf{x}}^{(i)}$ $=C(t,\textbf{x}^{(i)},\textbf{k}^{(i)}).$ (2)
For each training step, $\textbf{k}^{(i)}$ values are updated for each patient, $i$, using gradient descent optimization.
For each time step within the forecast horizon, $H$, forecasts, $\hat{\textbf{y}}$, are generated by the model, $f_{\theta}$, as a function of static features, s, treatment concentration curves, $\widecheck{\textbf{x}}$, and historical glucose values y recorded in the time prior to $t$ based on the lag time, $L$,
$\hat{\textbf{y}}^{(i)}_{t:t+H}$ $=f_{\theta}(\textbf{y}^{(i)}_{t-L:t},\widecheck{\textbf{x}}^{(i)}_{t-L:t},\textbf{s}^{(i)}).$ (3)
In the proposed context, s includes patient age, weight, one-hot encoded subject identification number, and insulin pump type. $\widecheck{\textbf{x}}_{t-L:t}$ includes concentration curves for basal insulin, $\widecheck{\textbf{x}}_{\text{basal}}$, bolus insulin, $\widecheck{\textbf{x}}_{\text{bolus}}$, and CHO, $\widecheck{\textbf{x}}_{\text{CHO}}$. We include plasma treatment effects of CHO because, similar to drugs, it undergoes absorption and metabolism processes, resulting in glucose and other sugars entering the bloodstream.
The PK encoder training framework is outlined in Appendix A.6, which can be integrated into any deep learning architecture capable of handling exogenous variables alongside modeling the target variable.
3.2 Modeling Multiple Dose Effects
Multiple insulin doses may occur at close intervals in an event often referred to as “insulin stacking” as shown in Fig. 11. To model PK effects for multiple doses in an input window, we generate a concentration curve for each dose event and sum concentration curves across the windows. Under the assumption of linear pharmacokinetics (outlined in 1.1 and Appendix B), we show that the overall effect, or concentration, of $n$ doses of the same drug is equal to the sum of the individual effects which is verified mathematically in Proposition A.2 in Appendix A.5. To perform vectorized operations that enable generating concentration curves for multiple doses in an input window, the PK encoder leverages a matrix, $\textbf{w}\in\mathbb{R}^{L\times L}$, where $L$ refers to the lag time. This approach allows for efficient modeling of multiple effects in $\mathcal{O}(1)$ time. Each row of w represents a time count up to the input time where the count shifts one time step for each subsequent row. Concentration curves are generated for each row of w using the dose at the corresponding time index. If no dose is administered at a particular time, $t$, the corresponding row of w indexed by $t$ will be zeroed out. The final concentration curve feature, which can reflect the impact of multiple doses, is generated by aggregating the individual concentration curves across time steps:
$\widecheck{\textbf{x}}^{(i)}$ $=\sum_{t=0}^{L}C(\textbf{w}_{t,j}\cdot f,\textbf{x}^{(i)}_{t},\textbf{k}^{(i)})$ (5)
Here, $f$ is the frequency of the data, $\textbf{x}^{(i)}\in\mathbb{R}^{L\times 1}$ a sparse exogenous time series of treatment, and $\widecheck{\textbf{x}}^{(i)}\in\mathbb{R}^{1\times L}$ is the new concentration curve feature for patient, $i$.
Figure 4: Our hybrid global-local architecture combines global model parameters shared across patients with patient-specific pharmacokinetic (PK) parameters. Sparse treatment time series features, such as medication doses, are input into the PK encoder to generate patient-specific treatment effect curves, $\check{\textbf{x}}$. PK encoder output, $\check{\textbf{x}}$, and any other exogenous features, x, are used as deep learning model inputs. During each training step, the model learns relevant hidden states and output parameters, $\hat{\theta}$, across patients, while $\textbf{k}^{(i)}$ values are updated for each patient, $i$, using gradient descent optimization.
3.3 Modeling Dose-Dependent Effects
A majority of drugs are assumed to adhere to linear PK, where it is assumed that there is a linear relationship between drug dose and concentration in the body (Rosenbaum, 2017). A characteristic of this assumption is that PK parameters are constant and do not change with dose. In practice, a drug’s PK parameters can vary depending on factors such as the drug’s properties, the type of injection, and the injection site. Nonlinear PK occurs when the relationship between drug dose and body concentration is not linear, leading to changes in PK parameters with dose. For instance, subcutaneous injections are administered into the subcutaneous tissue, where higher doses can saturate elimination pathways, causing slower absorption rates (Søeborg et al., 2009; Gradel et al., 2018). This variability in PK parameters can result in dose-dependent plasma drug concentration profiles.
CGMs administer the same type of insulin (rapid-acting) for basal and bolus doses. However, bolus insulin doses are generally much larger than basal doses as highlighted in Figs. 9 and 10 in Appendix A.1. To account for potential dose-dependent effects, we treat each dose range ($\textbf{k}_{\text{bolus}}$ and $\textbf{k}_{\text{basal}}$) as a separate entity. By treating each dose range as a separate entity, we can capture potential PK variations without resorting to a fully nonlinear PK model. This method simplifies the modeling process by maintaining linear assumptions within each dose category. It avoids the complexity of a full nonlinear model while still accounting for potential differences between the two categories of low-dose (basal) and high-dose (bolus) insulin kinetics. Specifically, we estimate $\textbf{k}_{\text{basal}}$ and $\textbf{k}_{\text{bolus}}$ values. For a single dose of basal insulin, $\textbf{x}_{\text{basal}}$, or bolus insulin, $\textbf{x}_{\text{bolus}}$, administered at time, $t$, we obtain the plasma insulin concentration at that time for patient, $i$:
$\widecheck{\textbf{x}}^{(i)}_{\text{basal}}$ $=C(t,\textbf{x}^{(i)}_{\text{basal}},\textbf{k}_{\text{basal}}^{(i)}),$ (6)
$\widecheck{\textbf{x}}^{(i)}_{\text{bolus}}$ $=C(t,\textbf{x}^{(i)}_{\text{bolus}},\textbf{k}_{\text{bolus}}^{(i)}).$ (7)
Treating each dose as a different entity is an approximation and may not capture all nuances of dose-dependent kinetics. All physiological models are approximations of complex biological processes. Modeling dose-dependent effects in a more granular manner could potentially provide more precise results if such effects are present. However, our aim is to develop minimal models validated under established linear PK assumptions from the pharmacology literature, striking a balance between simplicity and accuracy. Our framework ensures broader applicability across various medications where a majority of drugs are assumed to adhere to linear PK (Rosenbaum, 2017).
The equations presented in the paper are shown in a patient-specific manner to highlight that the PK encoder generates individualized information. However, these equations can be applied to produce patient-specific outputs for multiple patients simultaneously using vectorized operations.
3.4 Models
To assess forecasting performance across various model architectures, we trained models with six different algorithms as baselines. We employed Exponential Smoothing (ETS; Hyndman 2008) as our statistical baseline model. Our deep learning baselines included the following models: Multi-Layer Perceptron (MLP; Rosenblatt 1958), NBEATS/NHITS Oreshkin et al. (2020); Challu et al. (2022), Recurrent Neural Network (RNN; Elman (1990); Cho et al. (2014)), Long-Short Term Memory (LSTM; Sak et al. (2014)), and Temporal Convolution Network with an MLP Decoder (TCN; Bai et al. (2018); van den Oord et al. (2016)). Implementations of these models were obtained from the Neuralforecast (Olivares et al., 2022) and StatsForecast (Garza et al., 2022) libraries. More information on these models and related prior work that leverages these are provided in Appendix A.2. Details on model training and hyperparameter selection are provided in Appendix A.3
3.5 Baselines
We compare our hybrid global-local architecture with the PK encoder to both local and global baselines:
Global Baseline Models.
Models were trained on all patient data to create a single global model:
  • Baseline. Models are trained with the univariate blood glucose signal to demonstrate the forecasting improvement from using exogenous variables, supporting our PK encoder.
  • Baseline Exogenous (Exog.). Models are trained with the univariate blood glucose signals, static features, and sparse exogenous features (CHO and insulin dose values).
  • Baseline Exogenous (Exog.) - ST. Models are trained with the univariate blood glucose signal, static features, and sparse exogenous features transformed via the sum total (ST) approach, which converts features into cumulative values across an input window (Rubin-Falcone et al., 2022).
Local Baseline Exogenous Models.
Patient-specific baseline exogenous models were compared against our hybrid global-local PK encoder model.
Local PK Encoder Models.
Patient-specific models using the PK encoder were compared against our hybrid global-local PK encoder model.
3.6 Evaluation
Trained models were used to generate rolling window forecasts at one-step intervals. Similar to prior work (Rubin-Falcone et al., 2022; Xie and Wang, 2020), we evaluate the blood glucose forecasts for each test set in the two datasets using two metrics: mean absolute error (MAE) and root mean squared error (RMSE) computed across 8 trials of individually trained models. Formulations for these metrics are provided in Appendix A.4. We evaluate model performance across all points within the specified time horizon. We also assess the model’s performance exclusively at time points where blood glucose levels reach or fall below critical thresholds relevant to hypoglycemic and hyperglycemic events (i.e., $\leq$70 mg/dL and $\geq$180 mg/dL). For the OhioT1DM dataset, timestamps with forward-filled values were omitted from evaluation.